# Warehouse storage and retrieval optimization via clustering, dynamic systems modeling, and GPU-accelerated routing

### v1

@@ 4.3. Experiment 3
Figure 5. Three snapshots of cluster distributions from experiment 3, illustrated by bivariate Gaussian probability density functions defined using each cluster’s mean and covariance matrix $\Sigma_{n}(\ell)$, $\ell=1,2,3$. Each panel shows a superposition of the corresponding probability density functions. The snapshots are taken at time $n=1$ (left), $n=10$ (middle), $n=100$ (right). Two picking orders now differ by approximately $19\%$ on average, and we only see a slight improvement in the separation and variance of the clusters as time progresses. A movie showing the evolution of the clusters for $n=1,\ldots,200$ can be found here: https://play.hb.se/media/clusters_random_perturbation/0_g0zx8ooj.
4.4. Quantitative comparison
As experiments 1, 2, and 3 show, the picking and resorting procedure leads in time to increased separation of the initial clusters, although this separation becomes less pronounced as the variability of the orders increases. To quantify this effect we measure the area of the triangle defined by the cluster center points, and see how this varies with time for the three different setups used in experiments 1, 2, and 3. In Figure 6 we show snapshots of the cluster centers for the clusters depicted in Figure 3 (taken at the same picking iterations $n=1,10,100$ as in Figure 3), and we can clearly see that the area is greatly increased with subsequent iterations. For experiment 1 the locations of the cluster centers vary fairly slowly, while for experiment 2 and especially 3 the locations vary more rapidly, and as a result the area fluctuates more for experiment 2 and 3. To increase signal strength we therefore average the calculations over 10 runs for each experiment type 1, 2, and 3, and compute the area as a function of time (number of iterations of the picking and resorting algorithm). The result is presented in Figure 7. For each experiment type, the average area increases from an initial average area of around $0.1$; for experiment 1 it has increased to around 6 after the first 20 iterations; for experiment 2 it has increased to around 3 after the first 20 iterations; for experiment 3 it has increased to around 1 after the first 20 iterations. In all cases the increase is immediately visible and of at least one order of magnitude in size.
Refer to caption222244446666888810101010222244446666888810101010
Refer to caption222244446666888810101010222244446666888810101010
Refer to caption222244446666888810101010222244446666888810101010
Figure 6. Cluster center points corresponding to the snapshots of the clusters in Figure 3 from experiment 1, with $x$ and $y$ coordinates on the axes. As in Figure 3, the snapshots are taken after $n=1$ (left), $n=10$ (middle), and $n=100$ (right) picking iterations. Movies showing how the location of the cluster centers change with every iteration $n=1,\ldots,200$ for the experiments depicted in Figures 3, 4, and 5 can be found here: https://play.hb.se/media/cluster_centers_unperturbed/0_5eassah3, https://play.hb.se/media/cluster_centers_static_perturbation/0_njd00ou2, and https://play.hb.se/media/cluster_centers_random_perturbation/0_lazlyssr.
Refer to caption010020000222244446666888810101010Time
Refer to caption010020000222244446666888810101010Time
Refer to caption010020000222244446666888810101010Time
@@ 4.5. Route optimization
Here we analyze the effects that the resorting algorithm has on route optimization, and showcase the approximate method of calculating optimal routes that we describe in Section 3. We do so by analyzing an arbitrary route of a customer order, consisting of 10 products with varying quantities. The order is presented in Table 1. This particular order is assumed to be recurring (as in Experiment 1 above), and at the initial stage, iteration 1, the articles are randomly distributed throughout the entire domain of the warehouse node positions. Since the customer order is clustered, it will trigger a reorganization of clustered products, causing them to relocate closer to the cluster center. As stated earlier, the clusters tend to become more concentrated with reorganization, suggesting that the picking area within each cluster decreases. At iteration 300 we should therefore see a decrease in the total picking distance for each customer order.
Table 1. Products and their quantities. The position of each article in the warehouse is encoded through a value between 00 and $99$ to describe its $(x,y)$ coordinate in terms of a flattened $10\times 10$ grid starting from bottom left, together with a value between 00 and $9$ to describe its $z$ coordinate.
@@
When evaluating all possible route permutations through a graph with $n$ vertices, the total number of unique paths is given by $n!$, assuming each vertex is visited exactly once. For each such permutation, $(n-1)$ transitions must be evaluated to compute the total route cost, resulting in $n!(n-1)$ operations in total. In the CUDA C implementation, each route permutation may be assigned to a separate thread, and within each thread, an internal loop is used to process the transitions between vertices. If $V$ denotes the number of vertices (i.e., the number of stops per route), and $N=n!$ is the total number of permutations, then the computational structure involves operations proportional to $V\times N$. This scaling becomes problematic as $n$ increases, since $n!(n-1)$ grows factorially and dominates memory consumption. Empirically, we observed that values of $N$ larger than 2,903,040 lead to memory errors in the CUDA kernel, likely due to this exponential increase. As a result, processing large values of $n$ without segmenting the computation or distributing it across multiple kernel launches becomes infeasible due to the limitations in available GPU memory, see Figure 8.
@@
Figure 8. Data from the Bellman Ford GPU version (RTX6000, 24GB VRAM) (left), and a comparison between $n!(n-1)$ and computational time (right).
@@
Hence, since the number of stop positions in our experiment is $n=10$, we use the approximate method for calculating an optimal route described in Section 3. Initially, the number of stops is 10 while the number of passed nodes is 27, and the order is picked over a large portion of the warehouse area. By iteration 300, the clustered order regions have shrunk, and the number of passed nodes is now 15, determined using the approximate method. At iteration 300, the reorganization has resulted in many articles being located at the same stopping node, reducing the number of combinations to $6!$. This reduction enables a full combinatorial calculation of the optimal route. The result of the simulation is shown in Figure 9. As can be seen in this case, the reorganization yields an 80% increase in capacity. The approximate method does not produce the exact same result as the optimal method, with an optimal distance of 12. However, given the challenges with large $n!$, the approximate method allows for results close to the optimal solution in real-time when using GPUs.
@@
Figure 9. Picking routes for the customer order from Table 1 at different iterations of the resorting algorithm. The stopping positions along the route are indicated by the larger nodes, and they are clustered into 3 color coded regions using the methodology from Section 3. At iteration 1 there are 10 stopping nodes, and the left panel shows the computed route. At iteration 300, the resorting algorithm has reduced the number of stopping positions to $6$, and the middle panel shows the computed approximately optimal route with length 15. For comparison, the right panel shows the optimal route at 300 iterations having length 12.

### v2

@@ 4.3. Experiment 3
Refer to captionRefer to captionRefer to caption
Figure 5. Three snapshots of cluster distributions from experiment 3, illustrated by bivariate Gaussian probability density functions defined using each cluster’s mean and covariance matrix $\Sigma_{n}(\ell)$, $\ell=1,2,3$. Each panel shows a superposition of the corresponding probability density function plots. The snapshots are taken at time $n=1$ (left), $n=10$ (middle), $n=100$ (right). Two picking orders now differ by approximately $19\%$ on average, and we only see a slight improvement in the separation and variance of the clusters as time progresses. A movie showing the evolution of the clusters for $n=1,\ldots,200$ can be found here: https://play.hb.se/media/clusters_random_perturbation/0_g0zx8ooj
4.4. Cluster separation
As experiments 1, 2, and 3 show, the picking and resorting procedure leads in time to increased separation of the initial clusters, although this separation becomes less pronounced as the variability of the orders increases. To quantify this effect we measure the area of the triangle defined by the cluster center points, and see how this varies with time for the three different setups used in experiments 1, 2, and 3. For experiment 1 the locations of the cluster centers vary fairly slowly, while for experiment 2 and especially 3 the locations vary more rapidly, and as a result the area fluctuates more for experiment 2 and 3. To increase signal strength we therefore average the calculations over 10 runs for each experiment type 1, 2, and 3, and compute the area as a function of time (number of iterations of the picking and resorting algorithm).
The result is presented in Figure 6. For all experiment types, the average area increases from an initial average area of around $0.1$; for experiment 1 it has increased to around 6 after the first 20 iterations; for experiment 2 it has increased to around 3 after the first 20 iterations; for experiment 3 it has increased to around 1 after the first 20 iterations. In all cases, the increase is immediate and exceeds one order of magnitude, providing quantitative evidence that the resorting procedure effectively improves cluster separation.
Refer to caption01002000224466881010IterationArea
Refer to caption01002000224466881010Iteration
Refer to caption01002000224466881010Iteration
@@ 4.4. Cluster separation
4.5. Quantitative analysis of cluster quality
Here we quantify and compare the increase in cluster concentration and separation for the three experiments using silhouette scores to evaluate cluster quality [30]. We first recall the definition of silhouette score, so let $i$ denote a data point assigned to a cluster $A$. Define
$a(i)=\frac{1}{|A|-1}\sum_{\begin{subarray}{c}j\in A\\ j\neq i\end{subarray}}d(i,j),$
as the average dissimilarity between $i$ and all other points in its own cluster $A$. Here $d(i,j)$ is the distance between the data points $i$ and $j$. For every other cluster $C\neq A$, compute the average dissimilarity of $i$ to all points in $C$, and let
$b(i)=\min_{C\neq A}\frac{1}{|C|}\sum_{j\in C}d(i,j).$
The silhouette value for point $i$ is then defined as
$s(i)=\frac{b(i)-a(i)}{\max\{a(i),\,b(i)\}},\qquad-1\leq s(i)\leq 1.$
The overall silhouette score of a clustering solution with $n$ points is obtained by averaging over all points:
$S=\frac{1}{n}\sum_{i=1}^{n}s(i),$
with values ranging from 1-1 to +1+1. Higher scores indicate more compact and well-separated clusters, while values near 0 indicate overlapping clusters and negative values suggest misclassification. As a rule of thumb, scores above about $0.5$ reflect strong structure, scores around 0.250.250.50.5 indicate moderate clustering, and values close to 0 or below denote weak or absent cluster structure.
As in the previous subsection we perform 10 runs for each experiment type 1, 2, and 3 (abbreviated Exp 1, Exp 2, Exp 3 below), and calculate the silhouette score at each iteration. Figure 7 shows the evolution of silhouette scores over 100 iterations for the three experiment types. Exp 1 increases rapidly and stabilizes around 0.45, Exp 2 reaches about 0.10, and Exp 3 increases to around zero. The confidence intervals separate clearly between the experiments.
Refer to captionSilhouette scoreIterationSilhouette trajectories with 95% CI
Figure 7. Silhouette score trajectories over 100 iterations for a warehouse with $10\times 10\times 10$ nodes. Solid lines show the mean silhouette score across runs, with shaded bands indicating 95% confidence intervals. Experiment 1 shows strong and stable improvement, experiment 2 reaches modest gains, and experiment 3 remains close to zero, indicating limited clustering performance under high noise.
To quantify these improvements, we compare final to initial silhouette scores within each run. Paired $t$-based confidence intervals indicate substantial improvement in Exp 1 ($\Delta=0.47$, 95% CI [0.41, 0.53]), smaller but positive improvement in Exp 2 ($\Delta=0.12$, 95% CI [0.09, 0.16]), and only marginal improvement in Exp 3 ($\Delta=0.06$, 95% CI [0.03, 0.08]). These results suggest that, for the parameter settings described in the beginning of Section 4, the method remains effective under low and, to some extent, moderate noise, but yields limited gains under high noise. This is in accordance with the visual evidence of Figures 36. We note that under these settings the pick list already covers a substantial fraction of warehouse article types, and this fraction grows further with added noise. (As we will see, when the ratio of pick-list to warehouse article types is smaller, the method performs well also under moderate to high noise, see Section 5.)
To assess whether the magnitude of improvement differs between experiments, we perform pairwise permutation tests on the run-level improvements (10 000 resamples). All comparisons are statistically significant after Bonferroni correction (Exp 1 vs. Exp 2: $p<0.0001$; Exp 1 vs. Exp 3: $p<0.0001$; Exp 2 vs. Exp 3: $p=0.0063$). Effect sizes are also large: Exp 1 dominates both Exp 2 and Exp 3 (Cohen’s $d\geq 5$, Cliff’s $\delta=1.00$ in both cases), while Exp 2 also outperforms Exp 3 with a large effect (Cohen’s $d=1.5$, Cliff’s $\delta=0.76$). For reference, Cohen’s $d$ values of 0.2, 0.5, and 0.8 are often interpreted as small, medium, and large effects, respectively [10]; for Cliff’s $\delta$, thresholds of 0.147, 0.33, and 0.474 have been suggested for small, medium, and large effects [8, 29]. With only three pairwise comparisons, the Bonferroni correction had minimal impact on the results and did not alter the overall conclusions. Thus, although performance degrades with noise, clustering improvements remain detectable even under moderate noise. The final values and statistical comparisons are summarized in Table 1.
Table 1. Silhouette score improvements ($\Delta$) with 95% confidence intervals, and pairwise statistical comparisons between experiments for a warehouse with $1000$ nodes. Pairwise $p$-values are from permutation tests (Bonferroni corrected). Effect sizes are reported as Cohen’s $d$ and Cliff’s $\delta$.
Exp 1 Exp 2 Exp 3
Mean improvement $\Delta$ 0.47 0.12 0.06
95% CI [0.41, 0.53] [0.09, 0.16] [0.03, 0.08]
Pairwise comparisons
Exp 1 vs Exp 2 $p<0.0001$, $d=5.0$, $\delta=1.0$
Exp 1 vs Exp 3 $p<0.0001$, $d=6.3$, $\delta=1.0$
Exp 2 vs Exp 3 $p=0.0063$, $d=1.5$, $\delta=0.76$
@@ 4.6. Route optimization
Here we analyze the effects that the resorting algorithm has on route optimization, and showcase the approximate method of calculating optimal routes that we describe in Section 3, performed over 10 independent runs. Each run starts from an arbitrary customer order, consisting of 10 products with varying quantities. The order is assumed to be recurring (as in Experiment 1 above), and at the initial stage, iteration 1, the articles are randomly distributed throughout the entire domain of the warehouse node positions. Since the customer order is clustered, it will trigger a reorganization of clustered products, causing them to relocate closer to the cluster center. As seen above, the clusters tend to become more concentrated with reorganization, suggesting that the picking area within each cluster decreases. At iteration 300 we should therefore see a decrease in the total picking distance for each customer order.
We first present one representative trial, choosing the case with the largest deviation from optimal among the 10 independent runs. In this instance, the obtained route length exceeded the optimal by 25%, which constitutes the worst performance observed under the tested settings. The order for this case is presented in Table 2.
Table 2. Products and their quantities. The position of each article in the warehouse is encoded through a value between 0 and $99$ to describe its $(x,y)$ coordinate in terms of a flattened $10\times 10$ grid starting from bottom left, together with a value between 0 and $9$ to describe its $z$ coordinate.
@@
When evaluating all possible route permutations through a graph with $n$ vertices, the total number of unique paths is given by $n!$, assuming each vertex is visited exactly once. For each such permutation, (n1)(n-1) transitions must be evaluated to compute the total route cost, resulting in n!(n1)n!(n-1) operations in total. In the CUDA C implementation (see Subsection 4.7 for details), each route permutation may be assigned to a separate thread, and within each thread, an internal loop is used to process the transitions between vertices. If $V$ denotes the number of vertices (i.e., the number of stops per route), and $N=n!$ is the total number of permutations, then the computational structure involves operations proportional to $V\times N$. This scaling becomes problematic as $n$ increases, since n!(n1)n!(n-1) grows factorially and dominates memory consumption. Empirically, we observed that values of $N$ larger than 2,903,040 lead to memory errors in the CUDA kernel, likely due to this exponential increase. As a result, processing large values of $n$ without segmenting the computation or distributing it across multiple kernel launches becomes infeasible due to the limitations in available GPU memory, see Figure 8.
@@
Figure 8. Data from the Bellman Ford GPU version (RTX6000, 24GB VRAM) (left), and a comparison between n!(n1)n!(n-1) and computational time (right).
@@
Hence, since the number of stop positions in our experiment is $n=10$, we use the approximate method for calculating an optimal route described in Section 3. Initially, the number of stops is 10 while the number of passed nodes is 27, and the order is picked over a large portion of the warehouse area. By iteration 300, the clustered order regions have shrunk, and the number of passed nodes is now 15, determined using the approximate method. This corresponds to a 44% reduction in route length at iteration 300. At this stage, the reorganization has resulted in many articles being located at the same stopping node, reducing the number of combinations to $6!$. This reduction enables a full combinatorial calculation of the optimal route. The result of the simulation is shown in Figure 9. The approximate method does not produce the exact same result as the optimal method, with an optimal distance of 12. However, given the challenges with large $n!$, the approximate method allows for results close to the optimal solution in real-time when using GPUs.
@@
Figure 9. Illustrative example of the clustering-based route optimization, showing picking routes for the customer order from Table 2 at different iterations of the resorting algorithm. This is the trial with the largest deviation from optimal among the 10 runs considered, where the obtained route length exceeds the optimal by 25% at final iteration. The stopping positions along the route are indicated by the larger nodes, and they are clustered into 3 color coded regions using the methodology from Section 3. At iteration 1 there are 10 stopping nodes, and the left panel shows the computed route. At iteration 300, the resorting algorithm has reduced the number of stopping positions to $6$, and the middle panel shows the computed approximately optimal route with length 15. For comparison, the right panel shows the optimal route after 300 iterations, with length 12.
While this example illustrates the worst-performing case, we next summarize results across all runs described in Table 3. The ratio of obtained to optimal route length had a mean of 1.08 (SD = 0.08, median = 1.09), ranging from 1.00 (exact optimal) to 1.25 (worst case). The 95% confidence interval for the mean ratio was [1.02, 1.14], indicating that the method typically produced solutions within 2–14% of the optimal route length despite requiring much less computation. No significant correlation was found between optimal route length and approximation ratio ($r=0.01$, $p=0.99$), indicating that performance was stable across different route lengths.
Table 3. Comparison between optimal route length and clustering-based route length across 10 selected cases. Diff = difference between obtained and optimal length. Ratio = obtained/optimal.
Case Optimal Obtained Diff Ratio
1 13 15 2 1.15
2 11 11 0 1.00
3 8 9 1 1.12
4 7 7 0 1.00
5 11 12 1 1.09
6 12 15 3 1.25
7 14 14 0 1.00
8 14 14 0 1.00
9 10 11 1 1.10
10 12 13 1 1.08
Mean 1.08
SD 0.08
95% CI [1.02, 1.14]
4.7. GPU Implementation
The Bellman–Ford algorithm was parallelized and implemented in CUDA C, targeting an NVIDIA RTX 6000 GPU (Turing architecture, 24 GB GDDR6). Each edge relaxation was assigned to one GPU thread, enabling full parallel coverage of all $E$ edges per iteration. The kernel was configured with 256 threads per block, resulting in $\lceil E/256\rceil$ blocks per launch. For a representative instance with $V=9716$ vertices and E=26  [^9] E=26\,999 directed edges, this configuration corresponds to 106 thread blocks per iteration. The graph data were stored in edge-list format using integer arrays u, v, and w for source vertices, destination vertices, and edge weights, respectively. The distance and predecessor arrays (dist, pred) were allocated in global memory as contiguous vectors of size $V\times N$, where $N$ denotes the number of source vertices processed simultaneously. This layout allows coalesced memory access and efficient indexing across threads. The total device memory footprint for all arrays was approximately 1 MB, which is negligible relative to the available VRAM, confirming that capacity was not a limiting factor.
Although the memory demand is small, performance is constrained by global memory bandwidth. Each thread performs repeated read–write operations to the shared distance arrays across all $N$ sources in every relaxation step. As $N$ increases, these concurrent accesses lead to high write contention and non-coalesced access patterns, saturating the GPU’s memory bus rather than its computational resources. This bottleneck is intrinsic to the Bellman–Ford algorithm when expressed in its fully parallel form: the very feature that enables massive edge-level concurrency also enforces global synchronization after every iteration. Any attempt to restructure the CUDA kernel to alleviate this effect would inherently reduce the degree of parallelism, negating the algorithm’s primary performance advantage. Execution times were measured using CUDA event timers. For the configuration above, the complete GPU execution—including (V1)(V-1) relaxation rounds and host synchronization—completed in approximately 100 ms, corresponding to a throughput of 2627 million traversed edges per second (MTEPS). Profiling confirmed that the kernel is memory-bound, with arithmetic utilization remaining below 20% of theoretical peak. All implementation details and datasets used for validation are available for reproducibility at https://github.com/mnbe1973/Bellman-Ford-Cuda-C.
Remark.
For validation, we verified that the CPU implementations of Dijkstra’s and Bellman–Ford algorithms produced identical shortest paths on representative test graphs before porting Bellman–Ford to CUDA C. This confirmed functional equivalence between the two approaches. Dijkstra’s algorithm, however, is not straightforward to express in a fully parallel GPU form due to its priority-queue dependence, whereas Bellman–Ford can relax all edges concurrently. Hence, the latter was chosen for GPU acceleration. Metaheuristic approaches such as ant colony optimization (ACO) were deemed outside the scope of this study. Preliminary CPU–GPU comparisons indicated numerical consistency, while performance gains align with previously reported GPU speedups for edge-relaxation-based algorithms.
5. Large-scale experiments
To assess the scalability of the proposed method, we perform an additional set of experiments on a warehouse configuration of size $100\times 100\times 10$, corresponding to 100 000100\,000 nodes. The parameter settings otherwise follow those of Section 4: each picking order consists of a group of 20 customer orders, each with 10 article types and 1–10 parcels of every such article type. This design preserves comparability with Experiments 1–3 above, while allowing us to examine performance in a substantially larger fulfillment center. In this section we thus have the following parameter settings:
  1. (1)
    The warehouse contains $100\times 100\times 10$ picking nodes located at $(i,j,k)\in[1,100]\times[1,100]\times[1,10]$. This means that $n_{x}=n_{y}=100$ and $n_{z}=10$ in (1).
  2. (2)
    The number of article types is N=89  [^0] N=89\,000, and we assume that the maximum number of articles at a node is $10$ for all article types, so $M_{a_{ijk}}=10$ for all $(i,j,k)$. Thus, the state space $X$ is
    $X=\{(a_{ijk},m_{ijk})_{ijk}:a_{ijk}\in[0,89\,000],\ m_{ijk}\in[0,10]\}.$
  3. (3)
    The initial warehouse state $x_{0}$ is randomized with the following constraints: It contains $1100$ xyxy-coordinate pairs $(i_{l},j_{l})$, $0\leq l\leq 1099$, such that the nodes $(i_{0},j_{0},k),\ldots,(i_{1099},j_{1099},k)$ are empty for all $k\in[1,10]$. This corresponds to $1100$ empty racks, for a total of 11 00011\,000 empty nodes. A unique article type $a_{ijk}\in[1,89\,000]$ is assigned to each of the remaining 89 00089\,000 nodes, and we assume that the initial warehouse state is fully stocked, so the balance at each nonempty node $10$.
  4. (4)
    An order $o$ to be picked is a collection of $20$ individual customer or purchase orders. A purchase order consists of $10$ article types with between $1$ and $10$ parcels of each type. The $20$ purchase orders are mutually disjoint, so $o$ consists of $200$ article types. In each picking iteration, the $20$ purchase orders are divided into three groups using $K$-means clustering based on where the articles are stored in the warehouse at that time (with $K=3$ reflecting the classical ABC demand classification scheme, as discussed in Section 1.1). Each such group is a cluster.
An important consequence of this scaling is that the ratio of picked article types to the total number of available article types is now much smaller than in the previous $10\times 10\times 10$ setup. This provides a test of how clustering performs when orders occupy only a small fraction of the storage capacity. Under these conditions, we observed that the method performed well across all noise levels, including Experiment 3, which previously showed only marginal improvement in the smaller warehouse setting.
Due to the larger problem size, we restrict attention in this section to quantitative evaluation using silhouette scores and corresponding statistical validation. In contrast to Section 4, we do not include visual illustrations of cluster compactness or separation, but instead report results directly through statistical measures of cluster quality. Justification for the choice of $100\times 100\times 10$ nodes is provided below.
Figure 10 and Table 4 summarize the results for the large warehouse configuration. Substantial improvements in silhouette score were observed in all experiments, with the largest gains in Exp 1 ($\Delta=0.86$, 95% CI [0.81, 0.90]), followed by Exp 2 ($\Delta=0.50$, 95% CI [0.45, 0.54]) and Exp 3 ($\Delta=0.18$, 95% CI [0.15, 0.21]). Pairwise permutation tests confirmed that all differences between experiments were highly significant ($p<0.0001$ after Bonferroni correction). Effect sizes were uniformly large (Cohen’s $d>5.5$, Cliff’s $\delta=1.0$), indicating clear separation between noise levels. These results show that the clustering method scales effectively to a warehouse with 100 000100\,000 nodes and maintains stable performance even under the highest tested noise setting.
Refer to captionSilhouette scoreIterationSilhouette trajectories with 95% CI
Figure 10. Silhouette score trajectories over 100 iterations for the large warehouse configuration ($100\times 100\times 10$ nodes). Solid lines show the mean silhouette score across runs, with shaded bands indicating 95% confidence intervals. All experiments show substantial improvements, with higher noise (Exp 3) leading to smaller gains compared to Exp 1 and Exp 2.
Table 4. Silhouette score improvements ($\Delta$) with 95% confidence intervals, and pairwise statistical comparisons between experiments for a large warehouse with 100 000100\,000 nodes. Pairwise $p$-values are from permutation tests (Bonferroni corrected). Effect sizes are reported as Cohen’s $d$ and Cliff’s $\delta$.
Exp 1 Exp 2 Exp 3
Mean improvement $\Delta$ 0.86 0.50 0.18
95% CI [0.81, 0.90] [0.45, 0.54] [0.15, 0.21]
Pairwise comparisons
Exp 1 vs Exp 2 $p<0.0001$, $d=5.6$, $\delta=1.0$
Exp 1 vs Exp 3 $p<0.0001$, $d=12.2$, $\delta=1.0$
Exp 2 vs Exp 3 $p<0.0001$, $d=5.6$, $\delta=1.0$
In academic benchmarking, Duque-Jaramillo et al. [16] document a distribution warehouse with two sections, 24 rows per section, 66 columns per row and seven levels per column, giving a total of 22 176 storage slots and overall dimensions of $72\,\text{m}\times 190\,\text{m}$. This is approximately 22% of the 100 000100\,000 nodes scale in our large instance and about 22 times the 1000 node scale in our small instance. Alqahtani [2] analyses a retail distribution centre handling 40 SKUs (32 in detailed analysis) within a mixed block-storage and selective-rack layout, equivalent to roughly 0.04% of our large instance and 4% of our small instance.
In industry benchmarking, Digi-Key’s Product Distribution Center expansion (PDCe) in Thief River Falls, Minnesota, is a 2.2 million ft2 ($\approx 204\,387\,\text{m}^{2}$) facility—one of the ten largest warehouses in North America—designed to enable nearly triple its previous daily average of 27 000 packages, with dual conveyor systems spanning more than 27 miles [13, 14]. While facility-level SKU counts are not public, its throughput capacity of approximately 81 000 packages/day allows for a conservative order-of-magnitude estimate. Assuming an average of 5–7 SKUs per package, typical for electronics distributors, this corresponds to roughly 486 000486\,000 SKU-picks per day. Although SKU-picks/day is not identical to unique SKUs stored, this approximate throughput represents about five times the SKU-level scale of our large-instance simulation.
These comparisons place our large-instance simulation within the lower order of magnitude of SKU or picking scales observed in major operational warehouses, while also highlighting heterogeneity in SKU density, throughput, and layout. Detailed SKU-level and throughput data for the largest operators, including Amazon, are not publicly disclosed at the facility level due to commercial confidentiality, a limitation also noted by Zhuang et al. [36]. Taken together, these benchmarks indicate that the scale of our simulated instance is representative of modern fulfillment operations while remaining computationally tractable for GPU-based optimization.